OpenStudy (anonymous):

Solve x2 - 4x = 12 by completing the square, show work please!!

OpenStudy (anonymous):

b is -4 half of b is -2 this answer, -2 times itself equals 4 and now we add 4 to both sides x^2 - 4x + 4 = 12 +4 Do you know the next step?

hero (hero):

meverett, I understand what you're trying to do here, but your approach is useless without first explaining the general method of approach for completing the square.

hero (hero):

You should either provide the complete solution, or if you're going to do it your way, provide the general method of approach

OpenStudy (anonymous):

I motivate not hate ...

hero (hero):

You assume something is "hate" ing on you.

OpenStudy (anonymous):

Just transfer the 12 to the left and you can "break" it into (x+2)(x-6)=0 From that it follows, x=-2 and x=6

hero (hero):

So in general, when solving quadratic equations using the complete the square method, you begin with a general form ax^2 + bx = c. Then you add (b/2)^2 to both sides, creating a perfect square on the left side. Once we have a perfect square on the left side, we put it in the form (x-b)^2 then solve for x algebraically. In this particular case, we have x^2 -4x = 12. So a = 1, b = -4 and c = 12 Since b = -4, (b/2)^2 = =(-4/2)^2 = (-2)^2 = 4 so b/2 = 4 Now we add 4 to both sides: x^2-4x +4 = 12 +4 Then put the left side in the completed square form: (x-2)^2 = 16 Now we can solve for x algebraically: x - 2 = +-sqrt{16} x = +4 + 2 x = -4 + 2 x = 6 x = -2

OpenStudy (anonymous):

Thank you for the detailed explanation!!!

hero (hero):

yw